# Majorana fermion induced power-law scaling in the violations of the Wiedemann-Franz law

### v2

@@ IV.2 Inelastic scattering with only phase relaxation
where $\epsilon\in[0,1]$, and $I$ is the $2\times 2$ scattering matrix. $\epsilon=1$ denotes maximal coupling, while $\epsilon\to 0$ denotes no coupling [9]. Similar to the setup in Fig. 1 (a), we can relate the charge and heat currents to the voltage and temperature difference using Landauer-Buttiker scattering theory. We set the voltage of the left terminal at zero and the temperature at $T$, while the right terminal is at voltage $V$ and temperature $T+\Delta T$. Using Eq. (11), we relate the charge and heat currents to the voltage and temperature biases in matrix form as,
(I↑, e1cI↑, h1cI↑, e2cI↑, h2cI↑, e3cI↑, h3cI↑, e4cI↑, h4c) = (L↑;e12;cVL↑;e13;cVL↑;e14;cVL↑;e12;cTL↑;e13;cTL↑;e14;cTL↑;h12;cVL↑;h13;cVL↑;h14;cVL↑;h12;cTL↑;h13;cTL↑;h14;cTL↑;e22;cVL↑;e23;cVL↑;e24;cVL↑;e22;cTL↑;e23;cTL↑;e24;cTL↑;h22;cVL↑;h23;cVL↑;h24;cVL↑;h22;cTL↑;h23;cTL↑;e24;cTL↑;e32;cVL↑;e33;cVL↑;e34;cVL↑;e32;cTL↑;e33;cTL↑;e34;cTL↑;h32;cVL↑;h33;cVL↑;h34;cVL↑;h32;cTL↑;h33;cTL↑;h34;cTL↑;e42;cVL↑;e43;cVL↑;e44;cVL↑↑;e42;cTL↑;e43;cTL↑;e44;cTL↑;h42;cVL↑;h43;cVL↑;h44;cVL↑;h42;cTL↑;h43;cTL↑;h44;cT) (VV3V3ΔRΔTΔT ),
(I↑, e1qI↑, h1qI↑, e2qI↑, h2qI↑, e3qI↑, h3qI↑, e4qI↑, h4q) = (L↑;e12;qVL↑;e13;qVL↑;e14;qVL↑;e12;qTL↑;e13;qTL↑;e14;qTL↑;h12;qVL↑;h13;qVL↑;h14;qVL↑;h12;qTL↑;h13;qTL↑;h14;qTL↑;e22;qVL↑;e23;qVL↑;e24;qVL↑;e22;qTL↑;e23;qTL↑;e24;qTL↑;h22;qVL↑;h23;qVL↑;h24;qVL↑;h22;qTL↑;h23;qTL↑;e24;qTL↑;e32;qVL↑;e33;qVL↑;e34;qVL↑;e32;qTL↑;e33;qTL↑;e34;qTL↑;h32;qVL↑;h33;qVL↑;h34;qVL↑;h32;qTL↑;h33;qTL↑;h34;qTL↑;e42;qVL↑;e43;qVL↑;e44;qVL↑↑;e42;qTL↑;e43;qTL↑;e44;qTL↑;h42;qVL↑;h43;qVL↑;h44;qVL↑;h42;qTL↑;h43;qTL↑;h44;qT) (VV3V3ΔRΔTΔT ),
(I↓, e1cI↓, h1cI↓, e2cI↓, h2cI↓, e3cI↓, h3cI↓, e4cI↓, h4c) = (L↓;e12;cVL↓;e13;cVL↓;e14;cVL↓;e12;cTL↓;e13;cTL↓;e14;cTL↓;h12;cVL↓;h13;cVL↓;h14;cVL↓;h12;cTL↓;h13;cTL↓;h14;cTL↓;e22;cVL↓;e23;cVL↓;e24;cVL↓;e22;cTL↓;e23;cTL↓;e24;cTL↓;h22;cVL↓;h23;cVL↓;h24;cVL↓;h22;cTL↓;h23;cTL↓;e24;cTL↓;e32;cVL↓;e33;cVL↓;e34;cVL↓;e32;cTL↓;e33;cTL↓;e34;cTL↓;h32;cVL↓;he33;cVL↓;he34;cVL↓;h32;cTL↓;h33;cTL↓;h34;cTL↓;e42;cVL↓;e43;cVL↓;e44;cVL↓;e42;cTL↓;e43;cTL↓;e44;cTL↓;h42;cVL↓;h43;cVL↓;h44;cVL↓;h42;cTL↓;h43;cTL↓;h44;cT) (VV3V3ΔRΔTΔT ),
(I↓, e1qI↓, h1qI↓, e2qI↓, h2qI↓, e3qI↓, h3qI↓, e4qI↓, h4q) = (L↓;e12;qVL↓;e13;qVL↓;e14;qVL↓;e12;qTL↓;e13;qTL↓;e14;qTL↓;h12;qVL↓;h13;qVL↓;h14;qVL↓;h12;qTL↓;h13;qTL↓;h14;qTL↓;e22;qVL↓;e23;qVL↓;e24;qVL↓;e22;qTL↓;e23;qTL↓;e24;qTL↓;h22;qVL↓;h23;qVL↓;h24;qVL↓;h22;qTL↓;h23;qTL↓;e24;qTL↓;e32;qVL↓;e33;qVL↓;e34;qVL↓;e32;qTL↓;e33;qTL↓;e34;qTL↓;h32;qVL↓;he33;qVL↓;he34;qVL↓;h32;qTL↓;h33;qTL↓;h34;qTL↓;e42;qVL↓;e43;qVL↓;e44;qVL↓;e42;qTL↓;e43;qTL↓;e44;qTL↓;h42;qVL↓;h43;qVL↓;h44;qVL↓;h42;qTL↓;h43;qTL↓;h44;qT) (VV3V3ΔRΔTΔT ),
@@
${\begin{pmatrix}I_{1c}^{\uparrow,e}\\ I_{1c}^{\uparrow,h}\\ I_{2c}^{\uparrow,e}\\ I_{2c}^{\uparrow,h}\\ I_{3c}^{\uparrow,e}\\ I_{3c}^{\uparrow,h}\\ I_{4c}^{\uparrow,e}\\ I_{4c}^{\uparrow,h}\end{pmatrix}=-\begin{pmatrix}G(\mathcal{T}^{\uparrow;e}_{12})&G(\mathcal{T}^{\uparrow;e}_{13})&G(\mathcal{T}^{\uparrow;e}_{14})&S(\mathcal{T}^{\uparrow;e}_{12})&S(\mathcal{T}^{\uparrow;e}_{13})&S(\mathcal{T}^{\uparrow;e}_{14})\\ G(\mathcal{T}^{\uparrow;h}_{12})&G(\mathcal{T}^{\uparrow;h}_{13})&G(\mathcal{T}^{\uparrow;h}_{14})&S(\mathcal{T}^{\uparrow;h}_{12})&S(\mathcal{T}^{\uparrow;h}_{13})&S(\mathcal{T}^{\uparrow;h}_{14})\\ G(\mathcal{T}^{\uparrow;e}_{22})&G(\mathcal{T}^{\uparrow;e}_{23})&G(\mathcal{T}^{\uparrow;e}_{24})&S(\mathcal{T}^{\uparrow;e}_{22})&S(\mathcal{T}^{\uparrow;e}_{23})&S(\mathcal{T}^{\uparrow;e}_{24})\\ G(\mathcal{T}^{\uparrow;h}_{22})&G(\mathcal{T}^{\uparrow;h}_{23})&G(\mathcal{T}^{\uparrow;h}_{24})&S(\mathcal{T}^{\uparrow;h}_{22})&S(\mathcal{T}^{\uparrow;h}_{23})&S(\mathcal{T}^{\uparrow;h}_{24})\\ G(\mathcal{T}^{\uparrow;e}_{32})&G(\mathcal{T}^{\uparrow;e}_{33})&G(\mathcal{T}^{\uparrow;e}_{34})&S(\mathcal{T}^{\uparrow;e}_{32})&S(\mathcal{T}^{\uparrow;e}_{33})&S(\mathcal{T}^{\uparrow;e}_{34})\\ G(\mathcal{T}^{\uparrow;h}_{32})&G(\mathcal{T}^{\uparrow;h}_{33})&G(\mathcal{T}^{\uparrow;h}_{34})&S(\mathcal{T}^{\uparrow;he}_{32})&S(\mathcal{T}^{\uparrow;h}_{33})&S(\mathcal{T}^{\uparrow;h}_{34})\\ G(\mathcal{T}^{\uparrow;e}_{42})&G(\mathcal{T}^{\uparrow;e}_{43})&G(\mathcal{T}^{\uparrow;e}_{44})&S(\mathcal{T}^{\uparrow;e}_{42})&S(\mathcal{T}^{\uparrow;e}_{43})&S(\mathcal{T}^{\uparrow;e}_{44})\\ G(\mathcal{T}^{\uparrow;h}_{42})&G(\mathcal{T}^{\uparrow;h}_{43})&G(\mathcal{T}^{\uparrow;h}_{44})&S(\mathcal{T}^{\uparrow;h}_{42})&S(\mathcal{T}^{\uparrow;h}_{43})&S(\mathcal{T}^{\uparrow;h}_{44})\end{pmatrix}\begin{pmatrix}V\\ V_{3}\\ V_{3}\\ \Delta_{R}\\ \Delta T\\ \Delta T\end{pmatrix}},$ (40)
@@
${\begin{pmatrix}I_{1q}^{\uparrow,e}\\ I_{1q}^{\uparrow,h}\\ I_{2q}^{\uparrow,e}\\ I_{2q}^{\uparrow,h}\\ I_{3q}^{\uparrow,e}\\ I_{3q}^{\uparrow,h}\\ I_{4q}^{\uparrow,e}\\ I_{4q}^{\uparrow,h}\end{pmatrix}=-\begin{pmatrix}TS(\mathcal{T}^{\uparrow;e}_{12})&TS(\mathcal{T}^{\uparrow;e}_{13})&TS(\mathcal{T}^{\uparrow;e}_{14})&L(\mathcal{T}^{\uparrow;e}_{12})&L(\mathcal{T}^{\uparrow;e}_{13})&L(\mathcal{T}^{\uparrow;e}_{14})\\ TS(\mathcal{T}^{\uparrow;h}_{12})&TS(\mathcal{T}^{\uparrow;h}_{13})&TS(\mathcal{T}^{\uparrow;h}_{14})&L(\mathcal{T}^{\uparrow;h}_{12})&L(\mathcal{T}^{\uparrow;h}_{13})&L(\mathcal{T}^{\uparrow;h}_{14})\\ TS(\mathcal{T}^{\uparrow;e}_{22})&TS(\mathcal{T}^{\uparrow;e}_{23})&TS(\mathcal{T}^{\uparrow;e}_{24})&L(\mathcal{T}^{\uparrow;e}_{22})&L(\mathcal{T}^{\uparrow;e}_{23})&L(\mathcal{T}^{\uparrow;e}_{24})\\ TS(\mathcal{T}^{\uparrow;h}_{22})&TS(\mathcal{T}^{\uparrow;h}_{23})&TS(\mathcal{T}^{\uparrow;h}_{24})&L(\mathcal{T}^{\uparrow;h}_{22})&L(\mathcal{T}^{\uparrow;h}_{23})&L(\mathcal{T}^{\uparrow;h}_{24})\\ TS(\mathcal{T}^{\uparrow;e}_{32})&TS(\mathcal{T}^{\uparrow;e}_{33})&TS(\mathcal{T}^{\uparrow;e}_{34})&L(\mathcal{T}^{\uparrow;e}_{32})&L(\mathcal{T}^{\uparrow;e}_{33})&L(\mathcal{T}^{\uparrow;e}_{34})\\ TS(\mathcal{T}^{\uparrow;h}_{32})&TS(\mathcal{T}^{\uparrow;h}_{33})&TS(\mathcal{T}^{\uparrow;h}_{34})&L(\mathcal{T}^{\uparrow;he}_{32})&L(\mathcal{T}^{\uparrow;h}_{33})&L(\mathcal{T}^{\uparrow;h}_{34})\\ TS(\mathcal{T}^{\uparrow;e}_{42})&TS(\mathcal{T}^{\uparrow;e}_{43})&TS(\mathcal{T}^{\uparrow;e}_{44})&L(\mathcal{T}^{\uparrow;e}_{42})&L(\mathcal{T}^{\uparrow;e}_{43})&L(\mathcal{T}^{\uparrow;e}_{44})\\ TS(\mathcal{T}^{\uparrow;h}_{42})&TS(\mathcal{T}^{\uparrow;h}_{43})&TS(\mathcal{T}^{\uparrow;h}_{44})&L(\mathcal{T}^{\uparrow;h}_{42})&L(\mathcal{T}^{\uparrow;h}_{43})&L(\mathcal{T}^{\uparrow;h}_{44})\end{pmatrix}\begin{pmatrix}V\\ V_{3}\\ V_{3}\\ \Delta_{R}\\ \Delta T\\ \Delta T\end{pmatrix}},$ (41)

### v4

@@ Appendix B Expression for charge and heat current and Onsager coefficients with phase relaxation
Similar to the phase and momentum relaxation case, we can relate the charge and heat currents to the voltage and temperature difference using Landauer-Buttiker scattering theory. We set the voltage of the left terminal at zero and the temperature at $T$, while the right terminal is at voltage $V$ and temperature $T+\Delta T$. Using Eq. (11), we relate the charge and heat currents to the voltage and temperature biases in matrix form as,
(I↑, e1cI↑, h1cI↑, e2cI↑, h2cI↑, e3cI↑, h3cI↑, e4cI↑, h4c) = (L↑;e12;cVL↑;e13;cVL↑;e14;cVL↑;e12;cTL↑;e13;cTL↑;e14;cTL↑;h12;cVL↑;h13;cVL↑;h14;cVL↑;h12;cTL↑;h13;cTL↑;h14;cTL↑;e22;cVL↑;e23;cVL↑;e24;cVL↑;e22;cTL↑;e23;cTL↑;e24;cTL↑;h22;cVL↑;h23;cVL↑;h24;cVL↑;h22;cTL↑;h23;cTL↑;e24;cTL↑;e32;cVL↑;e33;cVL↑;e34;cVL↑;e32;cTL↑;e33;cTL↑;e34;cTL↑;h32;cVL↑;h33;cVL↑;h34;cVL↑;h32;cTL↑;h33;cTL↑;h34;cTL↑;e42;cVL↑;e43;cVL↑;e44;cVL↑↑;e42;cTL↑;e43;cTL↑;e44;cTL↑;h42;cVL↑;h43;cVL↑;h44;cVL↑;h42;cTL↑;h43;cTL↑;h44;cT) (VV3V3ΔTRΔTΔT ),
(I↑, e1qI↑, h1qI↑, e2qI↑, h2qI↑, e3qI↑, h3qI↑, e4qI↑, h4q) = (L↑;e12;qVL↑;e13;qVL↑;e14;qVL↑;e12;qTL↑;e13;qTL↑;e14;qTL↑;h12;qVL↑;h13;qVL↑;h14;qVL↑;h12;qTL↑;h13;qTL↑;h14;qTL↑;e22;qVL↑;e23;qVL↑;e24;qVL↑;e22;qTL↑;e23;qTL↑;e24;qTL↑;h22;qVL↑;h23;qVL↑;h24;qVL↑;h22;qTL↑;h23;qTL↑;e24;qTL↑;e32;qVL↑;e33;qVL↑;e34;qVL↑;e32;qTL↑;e33;qTL↑;e34;qTL↑;h32;qVL↑;h33;qVL↑;h34;qVL↑;h32;qTL↑;h33;qTL↑;h34;qTL↑;e42;qVL↑;e43;qVL↑;e44;qVL↑↑;e42;qTL↑;e43;qTL↑;e44;qTL↑;h42;qVL↑;h43;qVL↑;h44;qVL↑;h42;qTL↑;h43;qTL↑;h44;qT) (VV3V3ΔTRΔTΔT ),
(I↓, e1cI↓, h1cI↓, e2cI↓, h2cI↓, e3cI↓, h3cI↓, e4cI↓, h4c) = (L↓;e12;cVL↓;e13;cVL↓;e14;cVL↓;e12;cTL↓;e13;cTL↓;e14;cTL↓;h12;cVL↓;h13;cVL↓;h14;cVL↓;h12;cTL↓;h13;cTL↓;h14;cTL↓;e22;cVL↓;e23;cVL↓;e24;cVL↓;e22;cTL↓;e23;cTL↓;e24;cTL↓;h22;cVL↓;h23;cVL↓;h24;cVL↓;h22;cTL↓;h23;cTL↓;e24;cTL↓;e32;cVL↓;e33;cVL↓;e34;cVL↓;e32;cTL↓;e33;cTL↓;e34;cTL↓;h32;cVL↓;he33;cVL↓;he34;cVL↓;h32;cTL↓;h33;cTL↓;h34;cTL↓;e42;cVL↓;e43;cVL↓;e44;cVL↓;e42;cTL↓;e43;cTL↓;e44;cTL↓;h42;cVL↓;h43;cVL↓;h44;cVL↓;h42;cTL↓;h43;cTL↓;h44;cT) (VV3V3ΔTRΔTΔT ),
(I↓, e1qI↓, h1qI↓, e2qI↓, h2qI↓, e3qI↓, h3qI↓, e4qI↓, h4q) = (L↓;e12;qVL↓;e13;qVL↓;e14;qVL↓;e12;qTL↓;e13;qTL↓;e14;qTL↓;h12;qVL↓;h13;qVL↓;h14;qVL↓;h12;qTL↓;h13;qTL↓;h14;qTL↓;e22;qVL↓;e23;qVL↓;e24;qVL↓;e22;qTL↓;e23;qTL↓;e24;qTL↓;h22;qVL↓;h23;qVL↓;h24;qVL↓;h22;qTL↓;h23;qTL↓;e24;qTL↓;e32;qVL↓;e33;qVL↓;e34;qVL↓;e32;qTL↓;e33;qTL↓;e34;qTL↓;h32;qVL↓;he33;qVL↓;he34;qVL↓;h32;qTL↓;h33;qTL↓;h34;qTL↓;e42;qVL↓;e43;qVL↓;e44;qVL↓;e42;qTL↓;e43;qTL↓;e44;qTL↓;h42;qVL↓;h43;qVL↓;h44;qVL↓;h42;qTL↓;h43;qTL↓;h44;qT) (VV3V3ΔTRΔTΔT ),
@@
${\begin{pmatrix}I_{1c}^{\uparrow,e}\\ I_{1c}^{\uparrow,h}\\ I_{2c}^{\uparrow,e}\\ I_{2c}^{\uparrow,h}\\ I_{3c}^{\uparrow,e}\\ I_{3c}^{\uparrow,h}\\ I_{4c}^{\uparrow,e}\\ I_{4c}^{\uparrow,h}\end{pmatrix}=-\begin{pmatrix}G(\mathcal{T}^{\uparrow;e}_{12})&G(\mathcal{T}^{\uparrow;e}_{13})&G(\mathcal{T}^{\uparrow;e}_{14})&S(\mathcal{T}^{\uparrow;e}_{12})&S(\mathcal{T}^{\uparrow;e}_{13})&S(\mathcal{T}^{\uparrow;e}_{14})\\ G(\mathcal{T}^{\uparrow;h}_{12})&G(\mathcal{T}^{\uparrow;h}_{13})&G(\mathcal{T}^{\uparrow;h}_{14})&S(\mathcal{T}^{\uparrow;h}_{12})&S(\mathcal{T}^{\uparrow;h}_{13})&S(\mathcal{T}^{\uparrow;h}_{14})\\ G(\mathcal{T}^{\uparrow;e}_{22})&G(\mathcal{T}^{\uparrow;e}_{23})&G(\mathcal{T}^{\uparrow;e}_{24})&S(\mathcal{T}^{\uparrow;e}_{22})&S(\mathcal{T}^{\uparrow;e}_{23})&S(\mathcal{T}^{\uparrow;e}_{24})\\ G(\mathcal{T}^{\uparrow;h}_{22})&G(\mathcal{T}^{\uparrow;h}_{23})&G(\mathcal{T}^{\uparrow;h}_{24})&S(\mathcal{T}^{\uparrow;h}_{22})&S(\mathcal{T}^{\uparrow;h}_{23})&S(\mathcal{T}^{\uparrow;h}_{24})\\ G(\mathcal{T}^{\uparrow;e}_{32})&G(\mathcal{T}^{\uparrow;e}_{33})&G(\mathcal{T}^{\uparrow;e}_{34})&S(\mathcal{T}^{\uparrow;e}_{32})&S(\mathcal{T}^{\uparrow;e}_{33})&S(\mathcal{T}^{\uparrow;e}_{34})\\ G(\mathcal{T}^{\uparrow;h}_{32})&G(\mathcal{T}^{\uparrow;h}_{33})&G(\mathcal{T}^{\uparrow;h}_{34})&S(\mathcal{T}^{\uparrow;he}_{32})&S(\mathcal{T}^{\uparrow;h}_{33})&S(\mathcal{T}^{\uparrow;h}_{34})\\ G(\mathcal{T}^{\uparrow;e}_{42})&G(\mathcal{T}^{\uparrow;e}_{43})&G(\mathcal{T}^{\uparrow;e}_{44})&S(\mathcal{T}^{\uparrow;e}_{42})&S(\mathcal{T}^{\uparrow;e}_{43})&S(\mathcal{T}^{\uparrow;e}_{44})\\ G(\mathcal{T}^{\uparrow;h}_{42})&G(\mathcal{T}^{\uparrow;h}_{43})&G(\mathcal{T}^{\uparrow;h}_{44})&S(\mathcal{T}^{\uparrow;h}_{42})&S(\mathcal{T}^{\uparrow;h}_{43})&S(\mathcal{T}^{\uparrow;h}_{44})\end{pmatrix}\begin{pmatrix}V\\ V_{3}\\ V_{3}\\ \Delta T_{R}\\ \Delta T\\ \Delta T\end{pmatrix}},$ (47)
@@
${\begin{pmatrix}I_{1q}^{\uparrow,e}\\ I_{1q}^{\uparrow,h}\\ I_{2q}^{\uparrow,e}\\ I_{2q}^{\uparrow,h}\\ I_{3q}^{\uparrow,e}\\ I_{3q}^{\uparrow,h}\\ I_{4q}^{\uparrow,e}\\ I_{4q}^{\uparrow,h}\end{pmatrix}=-\begin{pmatrix}TS(\mathcal{T}^{\uparrow;e}_{12})&TS(\mathcal{T}^{\uparrow;e}_{13})&TS(\mathcal{T}^{\uparrow;e}_{14})&L(\mathcal{T}^{\uparrow;e}_{12})&L(\mathcal{T}^{\uparrow;e}_{13})&L(\mathcal{T}^{\uparrow;e}_{14})\\ TS(\mathcal{T}^{\uparrow;h}_{12})&TS(\mathcal{T}^{\uparrow;h}_{13})&TS(\mathcal{T}^{\uparrow;h}_{14})&L(\mathcal{T}^{\uparrow;h}_{12})&L(\mathcal{T}^{\uparrow;h}_{13})&L(\mathcal{T}^{\uparrow;h}_{14})\\ TS(\mathcal{T}^{\uparrow;e}_{22})&TS(\mathcal{T}^{\uparrow;e}_{23})&TS(\mathcal{T}^{\uparrow;e}_{24})&L(\mathcal{T}^{\uparrow;e}_{22})&L(\mathcal{T}^{\uparrow;e}_{23})&L(\mathcal{T}^{\uparrow;e}_{24})\\ TS(\mathcal{T}^{\uparrow;h}_{22})&TS(\mathcal{T}^{\uparrow;h}_{23})&TS(\mathcal{T}^{\uparrow;h}_{24})&L(\mathcal{T}^{\uparrow;h}_{22})&L(\mathcal{T}^{\uparrow;h}_{23})&L(\mathcal{T}^{\uparrow;h}_{24})\\ TS(\mathcal{T}^{\uparrow;e}_{32})&TS(\mathcal{T}^{\uparrow;e}_{33})&TS(\mathcal{T}^{\uparrow;e}_{34})&L(\mathcal{T}^{\uparrow;e}_{32})&L(\mathcal{T}^{\uparrow;e}_{33})&L(\mathcal{T}^{\uparrow;e}_{34})\\ TS(\mathcal{T}^{\uparrow;h}_{32})&TS(\mathcal{T}^{\uparrow;h}_{33})&TS(\mathcal{T}^{\uparrow;h}_{34})&L(\mathcal{T}^{\uparrow;he}_{32})&L(\mathcal{T}^{\uparrow;h}_{33})&L(\mathcal{T}^{\uparrow;h}_{34})\\ TS(\mathcal{T}^{\uparrow;e}_{42})&TS(\mathcal{T}^{\uparrow;e}_{43})&TS(\mathcal{T}^{\uparrow;e}_{44})&L(\mathcal{T}^{\uparrow;e}_{42})&L(\mathcal{T}^{\uparrow;e}_{43})&L(\mathcal{T}^{\uparrow;e}_{44})\\ TS(\mathcal{T}^{\uparrow;h}_{42})&TS(\mathcal{T}^{\uparrow;h}_{43})&TS(\mathcal{T}^{\uparrow;h}_{44})&L(\mathcal{T}^{\uparrow;h}_{42})&L(\mathcal{T}^{\uparrow;h}_{43})&L(\mathcal{T}^{\uparrow;h}_{44})\end{pmatrix}\begin{pmatrix}V\\ V_{3}\\ V_{3}\\ \Delta T_{R}\\ \Delta T\\ \Delta T\end{pmatrix}},$ (48)